OpenStudy (shamil98):

identify the curve by finding a Cartesian equation for the curve. r^2 cos 2theta = 1 so i'm guessing i use x^2 +y^2 = r^2 and r^2 = 1/cos2theta or sec 2theta how do i continue?

OpenStudy (zale101):

Use the right triangle reference. 

OpenStudy (shamil98):

yeah i thought of that where x^2 = r^2 cos^2 theta and y^2 = r^2 sin^2 theta but the 2theta kinda screws up the hope for reduction..unless i use a trig identity sec

OpenStudy (zale101):

I would use trig identity for that

OpenStudy (shamil98):

x^2 = cos^2 theta/ cos 2 theta = 1/1-tan^2 theta (after some simplifying) and then y^2 = 1/tan^2 theta - 1 so.. y^2 = -x^2 so x^2 + y^2 = 0

OpenStudy (shamil98):

i think that might be wrong actually, solutions manual says its supposed to be a hyperbola but isn't in hyperbola form ?.

OpenStudy (shamil98):

o

OpenStudy (zale101):

identity *

OpenStudy (shamil98):

i'm trying to see how i would get that in a hyperbola form (question asks to identify the curve by converting to cartesian )

OpenStudy (zale101):

Oh i seee!

OpenStudy (zale101):

r^2cos2theta=1 So do a trig identity to cos 2 theta \(cos2\theta=cos^2\theta-sin^2\theta\) plug that in \(r^2(cos^2\theta-sin^2\theta)=1 \) No distribute the r^2 in \(r^2cos^2\theta-r^2sin^2\theta=1 \) We know that the trig identity for \(rcos\theta=x\) and \(rsin\theta=y\)

OpenStudy (shamil98):

omg how did i not see that zale u genius

OpenStudy (zale101):

\(r^2cos^2\theta-r^2sin^2\theta=1\) \(x^2=(rcos\theta)^2\) \(y^2=(rsin\theta)^2\)

OpenStudy (shamil98):

ya i got it thnx

OpenStudy (zale101):

No problem.